# Latent Optimal Paths by Gumbel Propagation for Variational Bayesian Dynamic Programming

### v2

@@ Abstract
We propose the stochastic optimal path which solves the classical optimal path problem by a probability-softening solution. This unified approach transforms a wide range of DP problems into directed acyclic graphs in which all paths follow a Gibbs distribution. We show the equivalence of the Gibbs distribution to a message-passing algorithm by the properties of the Gumbel distribution and give all the ingredients required for variational Bayesian inference of a latent path, namely Bayesian dynamic programming (BDP). We demonstrate the usage of BDP in the latent space of variational autoencoders (VAEs) and propose the BDP-VAE which captures structured sparse optimal paths as latent variables. This enables end-to-end training for generative tasks in which models rely on unobserved structural information. At last, we validate the behaviour of our approach and showcase its applicability in two real-world applications: text-to-speech and singing voice synthesis. Our implementation code is avaliable at https://github.com/XinleiNIU/LatentOptimalPathsBayesianDP.
@@ 1 Introduction
Optimal paths are often required in many generative tasks such as speech, music, and language modeling (Kim et al., 2020; Li et al., 2022; Cai et al., 2019). These tasks involve the simultaneous identification of structured relationships between data and conditions. Obtaining optimal paths given a graph constraint with weights to achieve an improved fit based on extracted unobserved structural features. In the context of optimal path problems, such as finding the shortest path in a graph, dynamic programming (DP) efficiently computes the solution by breaking the problem down into several sub-problems and finding optimal solutions within the sub-problems iteratively.
@@
In this work, we explore a novel and unified method to obtain structured sparse optimal paths in the latent space of variational autoencoders (VAEs) with DP. Instead of a smoothed approximation for the classical optimal path problem, we propose the stochastic optimal path, which is a probabilistic softening solution by defining a Gibbs distribution where the energy function is the path score. We show this to be equivalent to a message-passing algorithm on the directed acyclic graphs (DAG) using the max and shift properties of the Gumbel distribution. To learn the latent optimal paths, we give tractable ingredients for variational Bayesian inference (i.e., likelihood and KL divergence) using DP, namely Bayesian dynamic programming (BDP), as well as an efficient sampling algorithm, which enables VAEs to obtain latent optimal paths within a DAG and achieve end-to-end training on generative tasks that rely on sparse unobserved structural relationships. We make the following contributions:
(1) We present a unified framework that gives a probabilistic softening of the classical optimal path problem on DAGs. We notably give efficient algorithms in linear time for sampling, computing the likelihood, and computing the KL divergence, thereby providing all of the ingredients required for variational Bayesian inference with a latent optimal path.
(2) We introduce BDP-VAE that learns sparse latent optimal paths. In the case of conditional generation, the data is not observed during inference, it is difficult to form the distribution statistics (i.e., edge weights) in the prior encoder. We give an alternative and flexible method to form the distribution statistics on the conditional prior by making use of a flow-based model.
@@ 2 Related Work
In many conditional generative tasks, models usually rely on structured dependencies of data and conditions, in which the dependencies are unobserved. Ren et al. (2020) and Liu et al. (2022) use a multiple training strategy by obtaining the sparse dependencies from an external DP-based techniques (McAuliffe et al., 2017) at first, then use the outputs as additional inputs to the model. Kim et al. (2020) integrates a DP on a Glow-based model to obtain unobserved monotonic alignment in parallel. Other models with structured latent representations such as HMMs (Rabiner, 1989) and PCFGs (Petrov & Klein, 2007) make strong assumptions about the model structure which could limit their flexibility and applicability. Instead of these works, we propose a unified framework to enable the VAEs (Kingma & Welling, 2014) to capture structural latent variables (i.e., sparse optimal paths), allowing for flexible adaptation to a variety of downstream tasks.
The attention mechanism is widely applied for obtaining unobserved dependencies in many seq-to-seq tasks. Deng et al. (2018) makes use of the properties of VAEs and captures the unobserved non-structural dependencies by learning latent alignment with attention in VAEs. However, to obtain structural constraints, attention strongly relies on model structures and other techniques such as DP to extract marginalization of the attention alignment distribution (Yu et al., 2016b, a). Different from latent alignment with attention, we target solving the optimal path problem in a unified framework that can be easily adapted to any structural constraint by defining DAGs. By capturing unobserved sparse shortest paths under the defined DAG in a latent space, our BDP-VAE facilitates the development of more explicit structural unobserved dependencies for a variety of applications. Secondly, attention mechanisms can be computationally expensive, especially for large input sequences. Solutions such as Chiu & Raffel (2018) reduce the computational complexity by involving a DP. Our method seeks stochastic optimal paths that occur in linear time with respect to edge numbers of DAGs (Corollary 4.11).
@@ 4.2 Gumbel Propagation
We now state the first main result in Lemma 4.4.
@@ 4.3 Sampling and Likelihood
Let paths $\mathbf{y}=(y_{1},y_{2},\dots,y_{\left|\mathbf{y}\right|})$ denotes the component of the random variable $Y$ defined in (10), given that $Y=(y_{1},y_{2},\dots,y_{\left|\mathbf{y}\right|})$. The probability of the transition $v\rightarrow u$ is
@@ 4.4 KL Divergence
Figure 1: A pipeline of BDP-VAE, BDP-VAE captures the unobserved sparse structural dependency (i.e., optimal paths) in the latent space in parallel training the model and allows gradient-based optimization for learning the edge weights $\mathbf{W}$.
@@ 5.2 Conditional Prior
We have provided a closed-form KL divergence within the family $\mathcal{D}(\mathcal{R},\cdot,\alpha)$ in Lemma 4.10 which may be convenient for unconditional generation by pre-setting the prior distribution statistics directly. In most conditional generation tasks, the non-accessible $\mathbf{x}$ during the inference phase in real applications leads to problems on the prior encoder when forming the edge weights $\mathbf{W}^{(0)}$ given information of $\mathbf{c}$ only, especially in the case that $\mathbf{x}$ has varying lengths $t$. To address this issue, we give a flexible solution for inferring feature information of $\mathbf{x}$ given $\mathbf{c}$ to form the edge weights $\mathbf{W}^{(0)}$ in the conditional prior. Inspired by Ma et al. (2019), we make use of a flow-based model to infer information of $\mathbf{x}$ condition on $\mathbf{c}$ and further form the $\mathbf{W}^{(0)}$. Assume there exists a series of invertible transformations random variables $\mathbf{x}$, such that
@@ 6 Experiments
Figure 2: Toy experiments on BDP to find stochastic optimal paths. The first row is a 5-node DAG and density plots with different $\alpha$. The second row is an 8-node DAG and density plots with different $\alpha$.
In this section, we conduct four experiments to verify our method from Section 4 and Section 5, and show its applicability on two real-world applications. To show the generalization of methods in Section 4, we extend BDP to two common application examples of computational graphs: monotonic alignment (MA) (Kim et al., 2020) and dynamic time warping (DTW)  (Sakoe & Chiba, 1978; Mensch & Blondel, 2018). Details of examples of computational graphs, pseudo-codes, and time complexity analysis are provided in Appendix E. We provide detailed model architecture used in our experiments in Appendix F and experimental details and setup in Appendix G.
@@ 6.2 Application: End-to-end Text-to-Speech
The task of TTS involves an unobserved monotonic hard alignment that maps phonemes to time intervals since the order of the phonemes should be preserved while the duration of each may vary in speech. Thus, TTS models generate speech according to corresponding phonemes in which the duration of each phoneme is discrete, structural, and unobserved (Mehrish et al., 2023). To show the BDP-VAE framework can be easily adapted into down-stream tasks, we redesign the popular non-end-to-end TTS model FastSpeech2 (Ren et al., 2020) into the BDP-VAE framework, namely BDPVAE-TTS, which can capture unobserved hard monotonic dependencies between phonemes and utterances jointly on both training and inference.
@@
We make a comparison with other end-to-end TTS models (Shen et al., 2018a; Kim et al., 2020; Lu et al., 2021) which target capturing phoneme dependencies. Shen et al. (2018a) models the unobserved temporal dependencies by an auto-regressive architecture with attention. Kim et al. (2020) integrate a monotonic alignment search in parallel in a Glow model (Kingma & Dhariwal, 2018) to obtain hard monotonic alignment. Lu et al. (2021) utilizes the latent Gaussian and captures soft monotonic alignment in the decoder by causality-masked self-attention. Among them, BDPVAE-TTS performs discrete alignment on both training and inference that ensures model consistency for training and inference. BDPVAE-TTS gets a better MCD and RTF than Shen et al. (2018a) and Kim et al. (2020), but gets higher MCD and RTF than Lu et al. (2021).
For the RTF, since BDPVAE-TTS involves a linear time consumption DP algorithm which causes higher inference time than VAENAR-TTS. We have discussed time complexity for each end-to-end TTS model in this experiment as below. However, the linear time consumption (Corollary 4.11) is the best we can do for solving a DP problem.
@@
For the MCD, even though BDPVAE-TTS obtains a higher MCD than VAENAR-TTS, BDPVAE-TTS captures discrete monotonic aligned paths in its latent space on training and inference phase which ensures model’s train and test consistency and improves model’s interpretability. VAENAR-TTS compresses the learned feature of spectrograms with conditions into Gaussian latent variables and uses these in its decoder for reconstruction. During the decoding process, these latent variables are used alongside phonemes to reconstruct the outputs. In contrast, the latent variables in BDPVAE-TTS are designed to capture phoneme-duration dependencies. This focus provides the decoder with a nuanced understanding of how the phoneme contributes to the overall speech utterance, which can be interpreted by the inference fundamental frequency (F0) in Figure 3. We provide detailed additional interpretations in Appendix H.
@@ 6.3 Application: End-to-end Singing Voice Synthesis
Two inference results are visualized in Figure 4 where red rectangles indicate that the temporal structure between the generated Mel-spectrogram and ground truth is almost identical. As expected, the decoder of BDP-VAE synthesizes spectrograms according to the extended conditions (i.e., the phonemes) mapping by the dependency of the latent monotonic path from the prior encoder. Thus the decoder synthesizes singing voice according to latent path-aligned phoneme conditions.
@@ 6.4 Verify Behaviour of Latent Optimal Path in BDP-VAE
To the best of our knowledge, there are no existing similar works that enable VAEs to obtain hard latent optimal paths given any defined DAG structures. Thus, inspired by the experimental designs in  Mensch & Blondel (2018) and  Van Den Oord et al. (2017), we set the a latent distribution $\mathcal{D}(\mathcal{R}_{\text{DTW}},\mathbf{W}=\{w\in\mathbf{W}\sim U(0,1)\},1)$ as the baseline due to lack of direct comparison. The random baseline still follows the DTW constraint with uniform edge weights, which can verify the behaviour of latent space in BPD-VAE.
@@ 7 Conclusion
We introduce a probabilistic softening solution to the classical optimal path problem on DAGs, as stochastic optimal paths. To achieve variational Bayesian inference with latent paths, we give efficient and tractable algorithms for likelihood and KL divergence within the family of path distributions in linear time with respect to edge numbers by dynamic programming with properties of the Gumbel distribution as message-passing, namely Bayesian dynamic programming with Gumbel propagation. We demonstrate the usage of stochastic optimal paths in the VAE framework by proposing BDP-VAE. BDP-VAE captures sparse optimal paths as latent representations given a DAG and further achieves end-to-end training for generative tasks that rely on unobserved structural relationships. We showed how the BDP finds latent optimal paths under general small DAGs and demonstrated the BDP-VAE with the computational graph of monotonic alignment on two real-world applications to achieve an end-to-end framework. We verified the behaviour and generalization of the latent optimal paths under the computational graph of dynamic time warping. We also studied the sensitivity of the hyper-parameter $\alpha$ and gave suggestions for real applications. Our experiments show the success of our approach on generative tasks where it achieves end-to-end training involving unobserved sparse structural optimal paths. Beyond VAE, the BDP can potentially be integrated into other probabilistic frameworks or planning in model-based reinforcement learning to obtain latent paths that leave for future extensions.
Limitations and Broader Impacts: As a discrete latent variable model, BDP-VAE uses the REINFORCE estimator which may lead to high gradient variance and slow convergence during training. This limitation may be solved by involving variance reduction techniques during training. This paper presents work whose goal is to advance the field of Machine Learning and applications of its downstream tasks. There are many potential societal consequences of our work, however, we do not foresee our methods bringing negative social impacts.

### v3

@@ Abstract
We propose the stochastic optimal path which solves the classical optimal path problem by a probability-softening solution. This unified approach transforms a wide range of DP problems into directed acyclic graphs in which all paths follow a Gibbs distribution. We show the equivalence of the Gibbs distribution to a message-passing algorithm by the properties of the Gumbel distribution and give all the ingredients required for variational Bayesian inference of a latent path, namely Bayesian dynamic programming (BDP). We demonstrate the usage of BDP in the latent space of variational autoencoders (VAEs) and propose the BDP-VAE which captures structured sparse optimal paths as latent variables. This enables end-to-end training for generative tasks in which models rely on unobserved structural information. At last, we validate the behavior of our approach and showcase its applicability in two real-world applications: text-to-speech and singing voice synthesis. Our implementation code is available at https://github.com/XinleiNIU/LatentOptimalPathsBayesianDP.
@@ 1 Introduction
Optimal paths are often required in many generative tasks such as speech, music, and language modelling (Kim et al., 2020; Li et al., 2022; Cai et al., 2019). These tasks involve the simultaneous identification of structured relationships between data and conditions. Finding optimal paths given a graph constraint with learned weights can highlight unobserved structural relationships that are not immediately apparent, thus potentially improving the interpretability and effectiveness of the models. In the context of optimal path problems, such as finding the shortest path in a graph, dynamic programming (DP) efficiently computes the solution by breaking the problem down into several sub-problems and finding optimal solutions within the sub-problems iteratively.
@@
In this work, we explore a novel and unified method to obtain structured sparse optimal paths with DP and showcase its application in the latent space of variational autoencoders (VAEs). Instead of a smoothed approximation for the classical optimal path problem, we propose the stochastic optimal path, which is a probabilistic softening solution by defining a Gibbs distribution where the energy function is the path score. We show this to be equivalent to a message-passing algorithm on the directed acyclic graphs (DAG) using the max and shift properties of the Gumbel distribution. To learn the latent optimal paths, we give tractable closed-form ingredients for variational Bayesian inference (i.e., likelihood and KL divergence) using DP, namely Bayesian dynamic programming (BDP), as well as an efficient sampling algorithm, which enables VAEs to obtain latent optimal paths within a DAG and achieve end-to-end training on generative tasks that rely on sparse unobserved structural relationships. We make the following contributions:
(1) We present a unified framework that gives a probabilistic softening of the classical optimal path problem on DAGs. We notably give efficient algorithms in linear time for sampling, computing the likelihood, and computing the KL divergence, thereby providing all the ingredients required for variational Bayesian inference with a latent optimal path.
(2) We introduce BDP-VAE framework that learns sparse optimal paths in the latent space. In the case of conditional generation, the data is not observed during inference, it is difficult to form the distribution statistics (i.e., edge weights) in the prior encoder. We give an alternative and flexible method to form the distribution statistics on the conditional prior by making use of a flow-based model.
@@ 2 Related Work
In many conditional generative tasks, models usually rely on structured dependencies of data and conditions, in which the dependencies are unobserved. Ren et al. (2020) and Liu et al. (2022) use a multiple training strategy by obtaining the sparse dependencies from an external DP-based techniques (McAuliffe et al., 2017) at first, then use the outputs as additional inputs to the model. Kim et al. (2020) integrates a DP on a Glow-based model to obtain unobserved monotonic alignment in parallel. Other models with structured latent representations such as HMMs (Rabiner, 1989) and PCFGs (Petrov & Klein, 2007) make strong assumptions about the model structure, which could limit their flexibility and applicability. Instead of these works, we propose a unified framework to enable the VAEs (Kingma & Welling, 2014) to capture structural latent variables (i.e., sparse optimal paths), allowing for flexible adaptation to a variety of downstream tasks.
The attention mechanism (Vaswani et al., 2017) is widely applied for obtaining unobserved dependencies in many seq-to-seq tasks. Deng et al. (2018) makes use of the properties of VAEs and captures the unobserved non-structural dependencies by learning latent alignment with attention in VAEs. However, to obtain structural constraints, attention strongly relies on model structures and other techniques such as DP to extract marginalization of the attention alignment distribution (Yu et al., 2016b, a). Different from latent alignment with attention, we target solving the optimal path problem in a unified framework that can be easily adapted to any structural constraint by defining DAGs (e.g., structural alignment). By capturing unobserved sparse shortest paths under the defined DAG in a latent space, our BDP-VAE facilitates the development of more explicit structural unobserved dependencies for a variety of applications. Secondly, attention mechanisms can be computationally expensive, especially for large input sequences. Solutions such as Chiu & Raffel (2018) reduce the computational complexity by involving a DP. Our method seeks stochastic optimal paths that occur in linear time with respect to edge numbers of DAGs (Corollary 4.11).
@@ 4.2 Gumbel Propagation
The motivation for constructing $Q_{v}$ is $Q_{v}$ allows us to set up a recursion on the entire DAG $\mathcal{R}$. We now state the first main result in Lemma 4.4.
@@ 4.3 Sampling and Likelihood
Let paths $\mathbf{y}=(y_{1},y_{2},\dots,y_{\left|\mathbf{y}\right|})$ denote the component of the random variable $Y$ defined in (10), given that $Y=(y_{1},y_{2},\dots,y_{\left|\mathbf{y}\right|})$. The probability of the transition $v\rightarrow u$ is
@@ 4.4 KL Divergence
Figure 1: A pipeline of BDP-VAE. BDP-VAE captures the unobserved sparse structural dependency (i.e., optimal paths on a DAG) in the latent space in parallel training the model and allows gradient-based optimization for learning the edge weights $\mathbf{W}$.
@@ 5.2 Conditional Prior
We have provided a closed-form KL divergence within the distribution family $\mathcal{D}(\mathcal{R},\cdot,\alpha)$ in Lemma 4.10 which may be convenient for unconditional generation by pre-setting the prior distribution statistics directly or has tractable $\mathbf{W}^{(0)}$.
In most conditional generation tasks, the non-accessible $\mathbf{x}$ during the inference phase in real applications leads to problems on the prior encoder when forming the edge weights $\mathbf{W}^{(0)}$ given information on $\mathbf{c}$ only, especially in the case that $\mathbf{x}$ has varying lengths $t$. To address this issue, we give a flexible solution for inferring feature information of $\mathbf{x}$ given $\mathbf{c}$ to form the edge weights $\mathbf{W}^{(0)}$ in the conditional prior. Inspired by Ma et al. (2019), we make use of a flow-based model  [^2] The flow-based conditional prior strategy is proposed to solve the problem of forming edge weights $\mathbf{W}^{(0)}$ in case of $\mathbf{W}^{(0)}$ is intractable if $\mathbf{x}$ is non-accessible with varying lengths $t$. In practical applications, this strategy is not mandatory if the task has a known and fix prior $\mathcal{D}(\mathcal{R},\mathbf{W}^{0},\alpha)$ or tractable $\mathbf{W}^{(0)}$ (i.e., Lemma 4.10). as the conditional prior to infer information about $\mathbf{x}$ condition on $\mathbf{c}$ and further obtain edge weights $\mathbf{W}^{(0)}$.
Assume there exists a series of invertible transformations of random variables $\mathbf{x}$, such that
@@ 6 Experiments
Figure 2: Toy experiments on BDP to find stochastic optimal paths under randomly generated DAGs. The first row is a 5-node DAG and its density plots with different $\alpha$ value. The second row is an 8-node DAG and its density plots with different $\alpha$ value.
In this section, we conduct four experiments to verify our methods from Section 4 and Section 5, and show its applicability on two real-world applications. To show the generalization of methods in Section 4, we extend BDP to two common application examples of computational graphs: monotonic alignment (MA) (Kim et al., 2020) and dynamic time warping (DTW)  (Sakoe & Chiba, 1978; Mensch & Blondel, 2018). Details of examples of computational graphs, pseudo-codes, and time complexity analysis are provided in Appendix E. We provide detailed model architecture used in our experiments in Appendix F and experimental details and setup in Appendix G.
@@ 6.2 Application: End-to-end Text-to-Speech
The task of TTS involves an unobserved monotonic hard alignment that maps phonemes to time intervals, since the order of the phonemes should be preserved while the duration of each may vary in speech. Thus, TTS models generate speech according to corresponding phonemes in which the duration of each phoneme is discrete, structural, and unobserved (Mehrish et al., 2023). To show the BDP-VAE framework can be easily adapted into down-stream tasks, we redesigned the popular non-end-to-end TTS model FastSpeech2 (Ren et al., 2020) into the BDP-VAE framework, namely BDPVAE-TTS, which can capture unobserved hard monotonic dependencies between phonemes and utterances jointly on both training and inference.
@@
We make a comparison with other end-to-end TTS models (Shen et al., 2018a; Kim et al., 2020; Lu et al., 2021) which target capturing phoneme dependencies. Shen et al. (2018a) models the unobserved temporal dependencies by an auto-regressive architecture with attention. Kim et al. (2020) integrate a monotonic alignment search in parallel in a Glow model (Kingma & Dhariwal, 2018) to obtain hard monotonic alignment. Lu et al. (2021) utilizes the latent Gaussian and captures soft monotonic alignment in the decoder by causality-masked self-attention. Among them, BDPVAE-TTS performs discrete monotonic alignment on both training and inference that ensures model consistency for training and inference. BDPVAE-TTS gets a better MCD and RTF than Shen et al. (2018a) and Kim et al. (2020), but gets higher MCD and RTF than Lu et al. (2021).
For the RTF, since BDPVAE-TTS involves a linear time consumption DP algorithm which causes higher inference time than VAENAR-TTS. We have discussed time complexity for capturing phoneme-to-frame dependencies during training for each end-to-end TTS model in this experiment as below. However, the linear time consumption (Corollary 4.11) is the best we can do for solving a DP problem.
@@
For the MCD, even though BDPVAE-TTS obtains a higher MCD than VAENAR-TTS, BDPVAE-TTS captures discrete monotonic aligned paths in its latent space on training and inference phase which ensures model’s train and test consistency and improves the model’s interpretability. VAENAR-TTS compresses the learned feature of spectrograms with conditions into Gaussian latent variables and uses these in its decoder for reconstruction. During the decoding process, these latent variables are used alongside phonemes to reconstruct the outputs. In contrast, the latent variables in BDPVAE-TTS are designed to capture phoneme-duration dependencies. This focus provides the decoder with a nuanced understanding of how the phoneme contributes to the overall speech utterance, which can be interpreted by the inference fundamental frequency (F0) in Figure 3. We provide detailed additional interpretations in Appendix H.
@@ 6.3 Application: End-to-end Singing Voice Synthesis
Two inference results are visualized in Figure 4 where red rectangles indicate that the temporal structure between the generated Mel-spectrogram and ground truth is almost identical. As expected, the decoder of BDP-VAE synthesizes spectrograms according to the extended conditions (i.e., the phonemes) mapping by the dependency of the latent monotonic path from the prior encoder. Therefore, the decoder synthesizes singing voice according to latent path-aligned phoneme conditions.
@@ 6.4 Verify Behaviour of Latent Optimal Path in BDP-VAE
To the best of our knowledge, there are no existing similar works that enable VAEs to obtain hard latent optimal paths given any defined DAG structures. Thus, inspired by the experimental designs in  Mensch & Blondel (2018) and  Van Den Oord et al. (2017), we set a latent distribution $\mathcal{D}(\mathcal{R}_{\text{DTW}},\mathbf{W}=\{w\in\mathbf{W}\sim U(0,1)\},1)$ as the baseline due to lack of direct comparison. The random baseline still follows the DTW constraint with uniform edge weights, which can verify the behaviour of latent space in BPD-VAE.
@@ 7 Conclusion
We introduce a probabilistic softening solution to the classical optimal path problem on DAGs, as stochastic optimal paths. To achieve variational Bayesian inference with latent paths, we give efficient and tractable algorithms for sampling, likelihood and KL divergence within the family of path distributions in linear time with respect to edge numbers by dynamic programming with properties of the Gumbel distribution as message-passing, namely Bayesian dynamic programming with Gumbel propagation. We demonstrate the usage of stochastic optimal paths in VAE framework and propose BDP-VAE. BDP-VAE captures sparse optimal paths as latent representations given a DAG and further achieves end-to-end training for downstream generative tasks that rely on unobserved structural relationships. We showed how the BDP finds stochastic optimal paths under general small toy DAGs and demonstrated the BDP-VAE with the computational graph of monotonic alignment on two real-world applications to achieve an end-to-end framework. We verified the behaviour and generalization of the latent optimal paths under the computational graph of dynamic time warping. We also studied the sensitivity of the hyper-parameter $\alpha$ and gave suggestions for real applications. Our experiments show the success of our approach on generative tasks where it achieves end-to-end training involving unobserved sparse structural optimal paths. Beyond VAE, the BDP can potentially be integrated into other probabilistic frameworks or planning in model-based reinforcement learning to obtain latent paths that leave for future extensions.
Limitations: As a discrete latent variable model, BDP-VAE uses the REINFORCE estimator which may lead to high gradient variance and slow convergence during training. This limitation may be solved by involving variance reduction techniques during training.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning and applications of its downstream tasks. There are many potential societal consequences of our work; however, we do not foresee our methods bringing negative social impacts.